# Distributed Model Checking in Graphs Classes

### v1

@@ 3.1.1 First-Order Formulas for Forests of Bounded Depth
$(F,L)\models\textsf{type}_{\gamma_{\overline{v}},\delta_{\overline{v}}}(\overline{v}).$
Now let $I\subseteq\textsf{Type}(k,d,\Lambda)$ be the set of all possible lca-types of the shape $\textsf{type}_{\gamma_{\overline{v}},\delta_{\overline{v}}}$ that can be defined by picking a labeled forest $(F,L)\in\mathcal{F}_{d}$ and a tuple $\overline{v}\in\textsf{true}(\varphi,F,L)$. By construction, $\varphi^{\prime}(\overline{x})=\bigvee_{\psi\in I}\psi(\overline{x})$ is such that $\textsf{true}(\varphi,F,L)=\textsf{true}(\varphi^{\prime},F,L)$ for every $(F,L)\in\mathcal{F}_{d}[\Lambda]$. ∎
As an example, let us consider the following formula using labels in $\{0,1\}$:
@@
which holds when $x_{1}$ is adjacent to $x_{2}$, $x_{1}$ is labeled 00, and $x_{2}$ and $x_{3}$ are both labeled 00. This formula can be rewritten as the lca-reduced formula
$\tilde{\varphi}(x_{1},x_{2},x_{3})=\left(\bigvee_{i\in[0,d]}\xi^{1}_{i}(x_{1},x_{2},x_{3})\right)\vee\left(\bigvee_{i\in[0,d-1]}\xi_{i}^{2}(x_{1},x_{2},x_{3})\right)\vee\left(\bigvee_{i\in[0,d-1]}\xi_{i}^{3}(x_{1},x_{2},x_{3})\right)$
@@
$\xi^{1}_{i}(x_{1},x_{2},x_{3})=\bigg{(}$ $\textsf{lca}_{i}(x_{1},x_{1})\wedge\textsf{lca}_{i+1}(x_{2},x_{2})\wedge\textsf{lca}_{i+1}(x_{3},x_{3})\wedge\textsf{lca}_{i}(x_{1},x_{2})\wedge\textsf{lca}_{i}(x_{1},x_{3})\wedge\textsf{lca}_{i}(x_{2},x_{3})\wedge$
@@
$\xi^{2}_{i}(x_{1},x_{2},x_{3})=\bigg{(}$ $\textsf{lca}_{i+1}(x_{1},x_{1})\wedge\textsf{lca}_{i}(x_{2},x_{2})\wedge\textsf{lca}_{i+2}(x_{3},x_{3})\wedge\textsf{lca}_{i}(x_{1},x_{2})\wedge\textsf{lca}_{i+1}(x_{1},x_{3})\wedge\textsf{lca}_{i}(x_{2},x_{3})\wedge$
@@
$\xi^{3}_{i}(x_{1},x_{2},x_{3})=\bigg{(}$ $\textsf{lca}_{i+1}(x_{1},x_{1})\wedge\textsf{lca}_{i+2}(x_{2},x_{2})\wedge\textsf{lca}_{i}(x_{3},x_{3})\wedge\textsf{lca}_{i+1}(x_{1},x_{2})\wedge\textsf{lca}_{i}(x_{1},x_{3})\wedge\textsf{lca}_{i}(x_{2},x_{3})\wedge$
@@
Observe that $\tilde{\varphi}$ is the basic normal form of $\varphi$. Indeed, for every $i\in[0,d]$, $\xi_{i}^{1}$ is an lca-type formula, and, for every $i\in[0,d-1]$, $\xi_{i}^{2}$ and $\xi_{i}^{3}$ are lca-type formulas. For instance, $\xi_{i}^{1}=\textsf{type}_{\gamma,\delta}$ with
@@
In Figure 1 we depict $F_{\xi^{1}_{i}}$, $F_{\xi^{2}_{i}}$, and $F_{\xi^{3}_{i}}$. If a labeled forest $(F,L)$ with label set $\{0,1\}$ satisfies $\varphi$, then $F$ contains a forest isomorphic to $F_{\xi^{1}_{i}}$, $F_{\xi^{2}_{i}}$ or $F_{\xi^{3}_{i}}$, for some $i\in[0,d]$.
@@
3.1.2 Quantifier Elimination of lca-Reduced Formulas
In this section, we show that it is possible to eliminate existential quantifiers on lca-reduced formulas, by augmenting the set of labels.
Definition 1.
Let $\Lambda$ be a finite set of labels, and let $d$ be a positive integer. A formula $\varphi(\overline{x})\in\text{FO}[\Lambda]$ is said reducible on $\mathcal{F}_{d}[\Lambda]$ if there exists a set $\widehat{\Lambda}$, and an lca-reduced formula $\widehat{\varphi}(\overline{x})\in\text{FO}[\widehat{\Lambda}]$ such that $\widehat{\Lambda}$ and $\widehat{\varphi}$ depend only on $\Lambda$, $\varphi$ and $d$, and, for every $(F,L)\in\mathcal{F}_{d}[\Lambda]$, there exists a $\widehat{\Lambda}$-labeling $\widehat{L}$ of $F$ such that
$\textsf{true}(\varphi,F,L)=\textsf{true}(\widehat{\varphi},F,\widehat{L}).$
Each element in the triplet $(\widehat{\Lambda},\widehat{\varphi},\widehat{L})$ of Definition 1 is called the reduction on $\mathcal{F}_{d}[\Lambda]$ of $\Lambda$, $\varphi$, and $L$, respectively. Note that we care that $\widehat{\Lambda}$ and $\widehat{\varphi}$ depend only on $\Lambda$, $\varphi$ and $d$ because, in our certification scheme later in the paper, we want the prover to give $\widehat{\Lambda}$ and $\widehat{\varphi}$ to the nodes as constant-size certificates. The following lemma is at the core of quantifier elimination.
Lemma 4.
Let $\varphi(\overline{x})=\exists\overline{y}\;\zeta(\overline{x},\overline{y})$ be a formula in $\text{FO}[\Lambda]$ where $\zeta$ is an lca-reduced formula. It holds that $\varphi$ is reducible on $\mathcal{F}_{d}[\Lambda]$.
@@ 3.1.2 Quantifier Elimination of lca-Reduced Formulas
Without loss of generality, let us assume that $|x|=k\geq 1$. (In the case where $\varphi$ has no free-variables, we can create a formula $\varphi(z)=\varphi$ for a dummy variable $z$.) We only deal with the case where $|y|=1$, i.e., with the elimination of a single quantified variable. Indeed, for the case where $|y|>1$ we can iteratively eliminate the quantifiers one by one, until we obtain a quantifier-free (actually an lca-reduced) formula. Using Lemma 1 we can assume without loss of generality that $\zeta$ is lca-reduced. Moreover, thanks to Lemma 3, we can also assume that $\zeta(\overline{x},y)$ is expressed in the basic normal form. That is, there exist $I\subseteq\textsf{Type}(k+1,d,\Lambda)$ such that $\zeta(\overline{x},y)=\bigvee_{\psi\in I}\psi(\overline{x},y)$. Since the existential quantifier commutes with the disjunction, we get that
@@
Let us fix an arbitrary $(F,L)\in\mathcal{F}_{d}[\Lambda]$. Let us first assume that each formula $\psi\in I$ is reducible on $\mathcal{F}_{d}[\Lambda]$ (we show how to do this reduction later in the proof). Under this assumption, let us denote by $\widehat{\Lambda}^{\psi}$, $\widehat{\psi}$, and $\widehat{L}^{\psi}$ the resulting reductions. We can relabel every set $\Lambda^{\psi}$, for $\psi\in I$, so that the sets $\widehat{\Lambda}^{\psi}\smallsetminus\Lambda$ are pairwise disjoint. Let us then define:
@@
$\widehat{\Lambda}=\bigcup_{\psi\in I}\widehat{\Lambda}^{\psi},\;\widehat{\varphi}=\bigvee_{\psi\in I}\widehat{\psi},\text{ and, for every node $u\in F$, }\;\widehat{L}(u)=\bigcup_{\psi\in I}\widehat{L}^{\psi}(u).$
@@
We get that $\widehat{\Lambda}$, $\widehat{\varphi}$, and $\widehat{L}$ are reductions on $\mathcal{F}_{d}[\Lambda]$ of $\Lambda$, $\varphi$, and $L$, respectively. Therefore, it just remains to prove that every $\psi\in\textsf{Type}(k+1,d,\Lambda)$ is reducible.
@@
Let $\gamma$ and $\delta$ such that $\psi(\overline{x},y)=\textsf{type}_{\gamma,\delta}(\overline{x},y)\in\textsf{Type}(k+1,d,\Lambda)$, where $y$ is identified with variable numbered $k+1$ in the definitions of $\gamma$ and $\delta$. Without loss of generality, we can assume that $\delta(1,k+1)$ is maximum (this value exists because $\varphi$ has at least one free variable). Let
$h=\delta(1,k+1),\;h_{x}=\delta(1,1),\text{ and }h_{y}=\delta(k+1,k+1).$
@@
In other words, $x_{1}$ of at depth $h_{x}$, $y$ is of depth $h_{y}$, and the least common ancestor of $x_{1}$ and $y$ (if exists) is at depth $h$ in $F_{\psi}$. We consider three cases, depending on the value of $\delta(1,k+1)$, and on the relative values of $h_{y}$ and $h$.
Case 1: $\delta(1,k+1)\geq 0$ and $h_{y}=h$.
@@
In this case $y$ is an ancestor of $x_{1}$ in $F_{\psi}$. Let $(F,L)\in\mathcal{F}_{d}[\Lambda]$. A node $w$ at depth $h$ is called a candidate if $L(w)=\gamma(k+1)$. A node at depth $h_{x}$ having an ancestor that is a candidate is called a good node. Let $\widehat{\Lambda}^{\psi}=\Lambda\cup\{\textsf{good}\}$. We define a $\widehat{\Lambda}^{\psi}$-labeling $\widehat{L}^{\psi}$ of $F$ as follows:
@@
$\widehat{L}^{\psi}(u)=\left\{\begin{array}[]{ll}L(u)\cup\{\textsf{good}\}&\mbox{if $u$ is a good node, }\\ L(u)&\mbox{otherwise.}\end{array}\right.$
@@
Let us now construct an lca-reduced formula $\widehat{\psi}$ that is equivalent to $\psi$. Let us denote by $\widehat{\gamma}:[k]\rightarrow 2^{\Lambda}$ and $\widehat{\delta}:[k]\times[k]\rightarrow[-1,d]$ the restrictions of $\gamma$ and $\delta$ to $[k]$ (i.e., for each $i,j\in[k]$, $\widehat{\gamma}(i)=\gamma(i)$ and $\widehat{\delta}(i,j)=\delta(i,j)$). Let $\xi=\textsf{type}_{\widehat{\gamma},\widehat{\delta}}$. Since $h_{y}=h$, we have that $F_{\xi}=F_{\psi}$, and $L_{\xi}(\overline{u}_{\xi})=L_{\psi}(\overline{u}_{\psi})$. Now, given a tuple of vertices $\overline{v}=(v_{1},\dots,v_{k})$, we have that $(F,L)\models\exists y,\psi(\overline{v},y)$ if and only if the following conditions hold:
@@
  • for every $i\in[k]$, $L(v_{i})=\gamma(i)=\widehat{\gamma}(i)$, and
  • the ancestor of $v_{1}$ at level $h$ in $F[\overline{v}]$ is labeled $\gamma(k+1)$, i.e., $v_{1}$ is a good node.
From Lemma 2 we obtain that the first two conditions imply that $(F,L)\models\psi(\overline{v})$. The third condition can be expressed as $\textsf{good}\in L(v_{1})$. Therefore, $\widehat{\psi}$ is defined as the conjunction of the two previous checks, that is,
$\widehat{\psi}(\overline{x})=\textsf{type}_{\widehat{\gamma},\widehat{\delta}}(\overline{x})\land\textsf{lab}_{\textsf{good}}(x_{1}).$
Therefore, $\textsf{true}(\psi,F,L)=\textsf{true}(\widehat{\psi},F,\widehat{L}^{\psi})$ for every $\Lambda$-labeled forest $(F,L)$ of depth at most $d$.
Case 2: $\delta(1,k+1)\geq 0$ and $h_{y}>h$.
Let $(F,L)\in\mathcal{F}_{d}[\Lambda]$. In this case, a node $v$ at depth $h_{y}$ is called a candidate if $L(v)=\gamma(k+1)$. A node $v$ at depth greater than $h$ is called a good node if at least one of its descendants is a candidate node. Finally, a node $v$ at depth $h$ is called a pivot, and we denote $\kappa(v)$ the number of children of $v$ that are good nodes.
@@
Given a tuple $\overline{v}$ of $k$ vertices, checking whether $F[\overline{v}]$ is isomorphic to $F_{\xi}$ can be achieved by merely using formula $\xi$. Let $z$ be the pivot in $F[\overline{v}]$, that is, $z$ is the ancestor of $v_{1}$ at level $h$. Let $\kappa(z,\overline{v})$ be the set of good children of $z$ contained in $F[\overline{v}]$. To eliminate the existential quantifier in $\varphi(\overline{x})$, it is sufficient to check whether one can extend $F[\overline{v}]$ by plugging a path from $z$ to a candidate node not contained in $F[\overline{v}]$. This condition holds if $\kappa(z,\overline{v})<\kappa(z)$, and it is exactly what is checked when considering the formula $\widehat{\psi}$.
Let us define $\widehat{\Lambda}^{\psi}=\Lambda\cup[0,k]\cup\{\infty,\textsf{good}\}$, and let us consider the $\widehat{\Lambda}^{\psi}$-labeling $\widehat{L}^{\psi}$ of $F$ defined as follows. Initially, for every node $u$ of $F$, $\widehat{L}^{\psi}(u)=L(u)$. We then update the label $\widehat{L}^{\psi}(u)$ of every node $u$ as follows.
  • If $u$ is a node at depth $h_{x}$, then let us denote $v$ the ancestor of $u$ at depth $h$. The label $\ell(\kappa(v))$ is added to $\widehat{L}^{\psi}(u)$ where
    $\ell(\kappa(v))=\begin{cases}\kappa(v)&\text{ if }\kappa(v)\leq k\\ \infty&\text{otherwise}.\end{cases}$
  • If $u$ is a good node, the label good is added to $\widehat{L}^{\psi}(u)$.
Observe that the two conditions above are not necessarily exclusive, and it may be the case that both labels $\ell(\kappa(v))$ and good are added to the set of labels of a the same node $u$.
Let us suppose that there exists a is a lca-reduced $\widehat{\Lambda}^{\psi}$-formula $\alpha(\overline{x})$ that checks whether $\kappa(z)>\kappa(z,\overline{x})$ where $z$ is the ancestor at level $h$ of $x_{1}$. We can then define $\widehat{\psi}$ as the lca-reduced $\widehat{\Lambda}^{\psi}$-formula
@@
and obtain that $\textsf{true}(\psi,F,L)=\textsf{true}(\widehat{\psi},F,\widehat{L}^{\psi})$ for every $\Lambda$-labeled forest $(F,L)$ of depth at most $d$. So, it is sufficient to show that $\alpha(\overline{x})$ exists.
For this purpose, observe first that $\kappa(z,\overline{x})\leq k$ as the forest $F_{\xi}$ has degree at most $k$ (because it has at most $k$ leaves). Therefore, $\alpha(\overline{x})$ holds whenever $\kappa(z)>k$, which is equivalent to the predicate $\textsf{lab}_{\infty}(x_{1})$. If $\kappa(z)\leq k$ then the value of $\kappa(z)$ is stored in the label of $x_{1}$. Let us denote by $z_{1},\dots,z_{p}$ the $p\leq k$ children of $z$ in $F_{\psi}$. For every $i\in[p]$, let us denote by $X_{i}$ the set of variables in $\overline{x}$ that are descendants of $z_{i}$. For every $q\in[p]$, we also denote by $K(p,q)$ the set of all subsets of $[p]$ of at most $q$ elements. If $q=\kappa(z)$ and $t=\kappa(z,\overline{x})$, then checking $q>t$ can be done by checking the existence of set $S\in K(p,q-1)$ satisfying that, for every $i\in S$, $X_{i}$ contains a good node, and, for every $i\notin S$, $X_{i}$ does not contain a good node. Therefore, $\alpha(\overline{x})$ can be defined as:
$\alpha(\overline{x})=\textsf{lab}_{\infty}(x_{1})\lor\beta(\overline{x}),$

### v2

@@ 3.1.1 First-Order Formulas on Forests of Bounded Depth
$(F,\ell)\models\textsf{type}_{\gamma_{\overline{v}},\delta_{\overline{v}}}(\overline{v}).$
Now let $I\subseteq\textsf{Type}(k,d,\Lambda)$ be the set of all possible lca-types $\textsf{type}_{\gamma_{\overline{v}},\delta_{\overline{v}}}$ that can be defined from a labeled forest $(F,\ell)\in\mathcal{F}_{d}$, and a tuple of vertives $\overline{v}\in\textsf{true}(\varphi,F,\ell)$. By construction, the function
$\varphi^{\prime}(\overline{x})=\bigvee_{\psi\in I}\psi(\overline{x})$
satisfies $\textsf{true}(\varphi,F,\ell)=\textsf{true}(\varphi^{\prime},F,\ell)$ for every $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$. ∎
Example.
Let us consider the following formula using labels in $\{0,1\}$, hence the two predicates $\textsf{lab}_{0}$ and $\textsf{lab}_{1}$:
@@
which holds when $x_{1}$ is adjacent to $x_{2}$, $x_{1}$ is labeled $1$, and $x_{2}$ and $x_{3}$ are both labeled 00. This formula can be rewritten as the lca-reduced formula
$\tilde{\varphi}(x_{1},x_{2},x_{3})=\left(\bigvee_{i\in[0,d-2]}\xi^{1}_{i}(x_{1},x_{2},x_{3})\right)\vee\left(\bigvee_{i\in[0,d-3]}\xi_{i}^{2}(x_{1},x_{2},x_{3})\right)\vee\left(\bigvee_{i\in[0,d-3]}\xi_{i}^{3}(x_{1},x_{2},x_{3})\right)$
@@
$\xi^{1}_{i}(x_{1},x_{2},x_{3})=\bigg{(}$ $\textsf{lca}_{i}(x_{1},x_{1})\wedge\textsf{lca}_{i+1}(x_{2},x_{2})\wedge\textsf{lca}_{i+1}(x_{3},x_{3})\wedge$
$\textsf{lca}_{i}(x_{1},x_{2})\wedge\textsf{lca}_{i}(x_{1},x_{3})\wedge\textsf{lca}_{i}(x_{2},x_{3})\wedge$
@@
$\xi^{2}_{i}(x_{1},x_{2},x_{3})=\bigg{(}$ $\textsf{lca}_{i+1}(x_{1},x_{1})\wedge\textsf{lca}_{i}(x_{2},x_{2})\wedge\textsf{lca}_{i+2}(x_{3},x_{3})\wedge$
$\textsf{lca}_{i}(x_{1},x_{2})\wedge\textsf{lca}_{i+1}(x_{1},x_{3})\wedge\textsf{lca}_{i}(x_{2},x_{3})\wedge$
@@
$\xi^{3}_{i}(x_{1},x_{2},x_{3})=\bigg{(}$ $\textsf{lca}_{i+1}(x_{1},x_{1})\wedge\textsf{lca}_{i+2}(x_{2},x_{2})\wedge\textsf{lca}_{i}(x_{3},x_{3})\wedge$
$\textsf{lca}_{i+1}(x_{1},x_{2})\wedge\textsf{lca}_{i}(x_{1},x_{3})\wedge\textsf{lca}_{i}(x_{2},x_{3})\wedge$
@@
Observe that $\tilde{\varphi}$ is the basic normal form of $\varphi$. Indeed, for every $i\in[0,d-2]$, $\xi_{i}^{1}$ is an lca-type formula, and, for every $i\in[0,d-3]$, $\xi_{i}^{2}$ and $\xi_{i}^{3}$ are lca-type formulas. For instance, $\xi_{i}^{1}=\textsf{type}_{\gamma,\delta}$ with
@@
Figure 2 depicts $F_{\xi^{1}_{i}}$, $F_{\xi^{2}_{i}}$, and $F_{\xi^{3}_{i}}$. If a labeled forest $(F,\ell)$ with label set $\{0,1\}$ satisfies $\varphi$, then $F$ contains a forest isomorphic to $F_{\xi^{1}_{i}}$, $F_{\xi^{2}_{i}}$ or $F_{\xi^{3}_{i}}$, for some $i\in[0,d-2]$.
@@
3.1.2 Quantifier Elimination of lca-Expressed Formulas
In this section, we show that it is possible to eliminate existential quantifiers on lca-expressed formulas, by augmenting the set of labels.
Definition 3.5.
Let $\Lambda$ be a finite set of labels, and let $d$ be a positive integer. A formula $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$ is said reducible on $\mathcal{F}_{d}[\Lambda]$ if there exists a set $\widehat{\Lambda}$, and an lca-reduced formula $\widehat{\varphi}(\overline{x})\in\mathrm{FO}[\widehat{\Lambda}]$ such that (1) $\widehat{\Lambda}$ and $\widehat{\varphi}$ depend only on $\Lambda$, $\varphi$, and $d$, and (2) for every $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$, there exists a $\widehat{\Lambda}$-labeling $\widehat{\ell}$ of $F$ such that
$\textsf{true}(\varphi,F,\ell)=\textsf{true}(\widehat{\varphi},F,\widehat{\ell}).$
Each element in the triplet $(\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ of Definition 3.5 is called the reduction on $\mathcal{F}_{d}[\Lambda]$ of $\Lambda$, $\varphi$, and $\ell$, respectively. We define a similar notion for local formulas.
Definition 3.6.
Let $\Lambda$ be a finite set of labels, and let $d$ be a positive integer. A formula $\varphi(\overline{x})\in FO[\Lambda]$ is said $r$-local reducible on $\mathcal{F}_{d}[\Lambda]$ if there exists a set $\widehat{\Lambda}$, and an lca-reduced formula $\widehat{\varphi}(\overline{x})\in\mathrm{FO}[\widehat{\Lambda}]$ such that (1) $\widehat{\Lambda}$ and $\widehat{\varphi}$ depend only on $\Lambda$, $\varphi$, and $d$, and (2) for every $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$, there exists a $\widehat{\Lambda}$-labeling $\widehat{\ell}$ of $F$ such that
$\textsf{true}(\varphi,F,\ell)=\textsf{true}(\widehat{\varphi},F,\widehat{\ell})=\textsf{true}(\widehat{\varphi}\wedge\mathrm{ball}(\overline{x},x_{1},r),F,\widehat{\ell}).$
In Definition 3.6, one insists on the fact that the reduction $\widehat{\varphi}$ of $\varphi$ must be $r$-local. This is called a $r$-local reduction.
Remark.
The fact that both Definitions 3.5 and 3.6 require $\widehat{\Lambda}$ and $\widehat{\varphi}$ to depend only on $\Lambda$, $\varphi$ and $d$ is not crucial for deciding (local) formulas, but will be used for the design of our certification scheme later in the paper (the prover gives $\widehat{\Lambda}$ and $\widehat{\varphi}$ to the nodes as constant-size certificates).
The following lemma is at the core of quantifier elimination.
Lemma 3.7.
Every formula $\varphi(\overline{x})=\exists\overline{y}\;\zeta(\overline{x},\overline{y})$ in $\mathrm{FO}[\Lambda]$, where $\zeta$ is an lca-reduced formula, is reducible on $\mathcal{F}_{d}[\Lambda]$.
@@ 3.1.2 Quantifier Elimination of lca-Expressed Formulas
Without loss of generality, let us assume that $|x|=k\geq 1$. (In the case where $\varphi$ has no free-variables, we can create a formula $\varphi(z)=\varphi$ for a dummy variable $z$.) We only deal with the case where $|y|=1$, i.e., with the elimination of a single quantified variable. Indeed, for the case where $|y|>1$ we can iteratively eliminate the quantifiers one by one, until we obtain a quantifier-free (actually an lca-reduced) formula. Using Lemma 3.1 we can assume without loss of generality that $\zeta$ is lca-reduced. Moreover, thanks to Lemma 3.4, we can also assume that $\zeta(\overline{x},y)$ is expressed in the basic normal form. That is, there exist $I\subseteq\textsf{Type}(k+1,d,\Lambda)$ such that $\zeta(\overline{x},y)=\bigvee_{\psi\in I}\psi(\overline{x},y)$. Since the existential quantifier commutes with the disjunction, we get that
@@
Let us fix an arbitrary labeled graph $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$. Let us first assume that each formula $\psi\in I$ is reducible on $\mathcal{F}_{d}[\Lambda]$ (we show how to do this reduction later in the proof). Under this assumption, let us denote by $\widehat{\Lambda}^{\psi}$, $\widehat{\psi}$, and ^$\widehat{\ell}^{\psi}$ the resulting reductions. We can relabel every set $\Lambda^{\psi}$, for $\psi\in I$, so that the sets $\widehat{\Lambda}^{\psi}\smallsetminus\Lambda$ are pairwise disjoint. Let us then define:
@@
$\widehat{\Lambda}=\bigcup_{\psi\in I}\widehat{\Lambda}^{\psi},\;\widehat{\varphi}=\bigvee_{\psi\in I}\widehat{\psi},\text{ and, for every node $u\in F$, }\;\widehat{\ell}(u)=\bigcup_{\psi\in I}\widehat{\ell}^{\psi}(u).$
@@
We get that $\widehat{\Lambda}$, $\widehat{\varphi}$, and $\widehat{\ell}$ are reductions on $\mathcal{F}_{d}[\Lambda]$ of $\Lambda$, $\varphi$, and $\ell$, respectively. Therefore, it just remains to prove that every $\psi\in\textsf{Type}(k+1,d,\Lambda)$ is reducible.
@@
Let $\gamma$ and $\delta$ be such that $\psi(\overline{x},y)=\textsf{type}_{\gamma,\delta}(\overline{x},y)\in\textsf{Type}(k+1,d,\Lambda)$, where $y$ is identified with variable numbered $k+1$ in the definitions of $\gamma$ and $\delta$. Let $s\in[k]$ such that $\delta(s,k+1)$ is maximum (this value exists because $\varphi$ has at least one free variable). We define
$h=\delta(s,k+1),\;h_{s}=\delta(s,s),\text{ and }h_{y}=\delta(k+1,k+1).$
@@
In other words, $x_{s}$ is at depth $h_{s}$, $y$ is of depth $h_{y}$, and the least common ancestor of $x_{s}$ and $y$ (if exists) is at depth $h$ in $F_{\psi}$. We consider three cases, depending on the value of $\delta(s,k+1)$, and on the relative values of $h_{y}$ and $h$.
Case 1: $\delta(s,k+1)\geq 0$ and $h_{y}=h$.
@@
In this case $y$ is an ancestor of $x_{s}$ in $F_{\psi}$. Let $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$. A node $w$ at depth $h$ is called a candidate if $\ell(w)=\gamma(k+1)$. A node at depth $h_{s}$ having an ancestor that is a candidate is called a good node. Let $\widehat{\Lambda}^{\psi}=\Lambda\cup\{\textsf{good}\}$. We define a $\widehat{\Lambda}^{\psi}$-labeling ^$\widehat{\ell}^{\psi}$ of $F$ as follows:
@@
^$\widehat{\ell}^{\psi}(u)=\left\{\begin{array}[]{ll}\ell(u)\cup\{\textsf{good}\}&\mbox{if $u$ is a good node, }\\ \ell(u)&\mbox{otherwise.}\end{array}\right.$
@@
Let us now construct an lca-reduced formula $\widehat{\psi}$ that is equivalent to $\psi$. Let us denote by $\widehat{\gamma}:[k]\rightarrow 2^{\Lambda}$ and $\widehat{\delta}:[k]\times[k]\rightarrow[-1,d-1]$ the restrictions of $\gamma$ and $\delta$ to $[k]$ (i.e., for each $i,j\in[k]$, $\widehat{\gamma}(i)=\gamma(i)$ and $\widehat{\delta}(i,j)=\delta(i,j)$). Let $\xi=\textsf{type}_{\widehat{\gamma},\widehat{\delta}}$. Since $h_{y}=h$, we have that $F_{\xi}=F_{\psi}$, and $\ell_{\xi}(\overline{u}_{\xi})=\ell_{\psi}(\overline{u}_{\psi})$. Now, given a tuple of vertices $\overline{v}=(v_{1},\dots,v_{k})$, we have that $(F,\ell)\models\exists y,\psi(\overline{v},y)$ if and only if the following conditions hold:
@@
  • for every $i\in[k]$, $\ell(v_{i})=\gamma(i)=\widehat{\gamma}(i)$, and
  • the ancestor of $v_{s}$ at level $h$ in $F[\overline{v}]$ is labeled $\gamma(k+1)$, i.e., $v_{s}$ is a good node.
From Lemma 3.3 we obtain that the first two conditions imply that $(F,\ell)\models\psi(\overline{v})$. The third condition can be expressed as $\textsf{good}\in\ell(v_{1})$. Therefore, $\widehat{\psi}$ is defined as the conjunction of the two previous expressions, that is,
$\widehat{\psi}(\overline{x})=\textsf{type}_{\widehat{\gamma},\widehat{\delta}}(\overline{x})\land\textsf{lab}_{\textsf{good}}(x_{s}).$
Therefore, $\textsf{true}(\psi,F,\ell)=\textsf{true}(\widehat{\psi},F,\widehat{\ell}^{\psi})$ for every $\Lambda$-labeled forest $(F,\ell)$ of depth at most $d$.
Case 2: $\delta(s,k+1)\geq 0$ and $h_{y}>h$.
Let $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$. In this case, a node $v$ at depth $h_{y}$ is called a candidate if $\ell(v)=\gamma(k+1)$. A node $v$ at depth greater than $h$ is called a good node if at least one of its descendants is a candidate node. Finally, a node $v$ at depth $h$ is called a pivot, and we denote by $\kappa(v)$ the number of children of $v$ that are good nodes.
@@
Given a tuple $\overline{v}$ of $k$ vertices, checking whether $F[\overline{v}]$ is isomorphic to $F_{\xi}$ can be achieved by merely using formula $\xi$. Let $z$ be the pivot in $F[\overline{v}]$, that is, $z$ is the ancestor of $v_{s}$ at level $h$. Let $\kappa(z,\overline{v})$ be the set of good children of $z$ contained in $F[\overline{v}]$. To eliminate the existential quantifier in $\varphi(\overline{x})$, it is sufficient to check whether one can extend $F[\overline{v}]$ by plugging a path from $z$ to a candidate node not contained in $F[\overline{v}]$. This condition holds if $\kappa(z,\overline{v})<\kappa(z)$, and it is exactly what is checked when considering the formula $\widehat{\psi}$.
Let us define $\widehat{\Lambda}^{\psi}=\Lambda\cup[0,k+1]\cup\{\textsf{good}\}$, and let us consider the $\widehat{\Lambda}^{\psi}$-labeling ^$\widehat{\ell}^{\psi}$ of $F$ defined as follows. Initially, for every node $u$ of $F$, ^$\widehat{\ell}^{\psi}(u)=\ell(u)$. We then update the label ^$\widehat{\ell}^{\psi}(u)$ of every node $u$ as follows.
  • If $u$ is a node at depth $h_{s}$, then let us denote $v$ the ancestor of $u$ at depth $h$. The label $\lambda$ is added to ^$\widehat{\ell}^{\psi}(u)$ where
    $\lambda=\begin{cases}\kappa(v)&\text{ if }\kappa(v)\leq k\\ k+1&\text{otherwise}.\end{cases}$
  • If $u$ is a good node, the label good is added to ^$\widehat{\ell}^{\psi}(u)$.
Observe that the two conditions above are not necessarily exclusive, and it may be the case that both labels $\lambda$ and good are added to the set of labels of a the same node $u$.
Let us suppose that there exists an lca-reduced $\widehat{\Lambda}^{\psi}$-formula $\alpha(\overline{x})$ that checks whether $\kappa(z)>\kappa(z,\overline{x})$ where $z$ is the ancestor at level $h$ of $x_{s}$. We can then define $\widehat{\psi}$ as the lca-reduced $\widehat{\Lambda}^{\psi}$-formula